OpenStudy (anonymous):

A line in the Paschen series of hydrogen has a wavelength of 1010nm. From what state did the electron originate?

OpenStudy (anonymous):

We can use teh Rydberg formula for this \[\frac{1}{\lambda}=R\left[\frac{1}{(n_2)^2}-\frac{1}{(n_1)^2}\right]\]where \(\lambda\) is the wavelength of emitted (or absorbed) light, \(R\) is teh Rydberg constant (\(R=1.097373\times10^{7}\rm{ m}^{-1}\)), and \(n_1\) and \(n_2\) are the initial and final energy respectively levels of the transition. For the Paschen series of emission, we know that all transitions are to the n=3 state. we now have enough information to solve for \(n_1\), given that we are told that the emission line is 1010 nm. Plugging in the numbers and solving for \(n_1\) gives a value of \(n_1=6.919\) which we can round up to 7. Therefore the transition originates in the n = 7 level of the atom.

OpenStudy (anonymous):

Thanks @JonnyMcA I had the same ? :)